Assume that a sample is used to estimate a population mean μ . Find the margin of error M.E. that corresponds to a sample of size 7 with a mean of 49.5 and a standard deviation of 6.3 at a confidence level of 90%.
Answer #1

critical value, critical value, t(a/2,n-1) = t(0.1/2,7-1) = 1.943

ME = t(a/2,n-1)*(sd/sqrt(n))
ME = 1.943*(6.3/sqrt(7))
me = 4.626625318